P(x)=6x3+5x2-2x-1


1
Expert's answer
2021-12-15T00:51:31-0500
"P(x)=6x^3+5x^2-2x-1"

"=6x^3+6x^2-x^2-x-x-1"

"=6x^2(x+1)-x(x+1)-(x+1)"

"=(x+1)(6x^2-x-1)"

Let "6x^2-x-1=0."


"D=(-1)^2-4(6)(-1)=25"

"x=\dfrac{1\pm\sqrt{25}}{2(6)}=\dfrac{1\pm5}{12}"

"x_1=\dfrac{1-5}{12}=-\dfrac{1}{3}, x_2=\dfrac{1+5}{12}=\dfrac{1}{2}"

Then


"6x^2-x-1=6(x+\dfrac{1}{3})(x-\dfrac{1}{2})"

"=(3x+1)(2x-2)"

"P(x)=6x^3+5x^2-2x-1"

"=(x+1)(3x+1)(2x-2)"

Zeros: "-1, -\dfrac{1}{3}, \dfrac{1}{2}"